OpenStudy (anonymous):

The length of a rectangle is the sum of two whole numbers. The width of the rectangle is the difference of the same two whole numbers. Using these facts, write a verbal expression for the area of the rectangle.

OpenStudy (anonymous):

length = a + b width = a - b (a + b) (a - b) = a^2 - b^2

OpenStudy (anonymous):

Ahhh.... Okay. That makes sense. Thank you very much!